Find the equations of the normal to the ellipse x2 + 4y2 = 65 at the points (1, 4) and (7, 2). Find the equation of the straight line joining the intersection of the normals to the origin. 


1
Expert's answer
2021-08-17T17:23:02-0400
"x^2 + 4y^2 = 65"

Differentiate both sides with respect to "x"


"2x+8yy'=0"

"y'=-\dfrac{x}{4y}"

Then the slope of the normal to the ellipse at the point "(x_0, y_0)" is


"slope=m=\dfrac{4y_0}{x_0}"

The equation of the normal in point-slope form is


"y-y_0=\dfrac{4y_0}{x_0}(x-x_0)"

The equation of the normal in slope-intercept form is


"y=\dfrac{4y_0}{x_0}x-3y_0"

Point "(1,4)"


"y=\dfrac{4(4)}{1}x-3(4)"

"y=16x-12"

Point "(7,2)"

"y=\dfrac{4(2)}{7}x-3(2)"

"y=\dfrac{8}{7}x-6"


Point of intersection


"16x-12=\dfrac{8}{7}x-6"

"52x=21"

"x=\dfrac{21}{52}"

"y=16(\dfrac{21}{52})-12=-\dfrac{72}{13}"

Point "(\dfrac{21}{52}, -\dfrac{72}{13})"


The equation of the straight line joining the intersection of the normals to the origin. 


"y=kx"

Substitute


"-\dfrac{72}{13}=k(\dfrac{21}{52})"

"k=-\dfrac{96}{7}"

The equation of the straight line joining the intersection of the normals to the origin.


"y=-\dfrac{96}{7}x"